OpenStudy (anonymous):

Find the derivative of y witrh respect to X: y=-3x^2 e^(x+1)?

OpenStudy (amistre64):

y = -3x^2 e^(x+1) use product rule and substitution. y' = r'l + rl' ; l' = e^u Du = e^(x+1)(1) y' = -6x e^(x+1) + -3x^2 e^(x+1) simplfy as needed/wanted

OpenStudy (amistre64):

y' = -3x e^(x+1) (2 + x) if I kept track ot it all :)